OpenStudy (anonymous):

Can someone please help me... If y varies directly as x and inversely as the square of z, and if y= 144 when x=4 and z=3, find y when x=3 and z=4?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

y = kx/z^2 k = yz^2/x k = 144(9)/4 = 324 y = 324x/z^2

OpenStudy (amistre64):

y = 324(3)/16 = 60.75 if i see it right

OpenStudy (anonymous):

thank you so much!!

OpenStudy (amistre64):

if I was right; youre welcome :)